OpenStudy (anonymous):

Solve the following system equations. 3y=12 2x+y=-16 and some of my answers are (-4,-10) (-10,4) (4,10) (3,3)

OpenStudy (anonymous):

\[3y=12\]\[y=12\div3\] y=4 Then plug y into the other equation to find x

OpenStudy (anonymous):

That isn't helping

OpenStudy (anonymous):

okay well your first equation is 3y=12 correct?

OpenStudy (anonymous):

Its all together

OpenStudy (anonymous):

isnt it 3y=12 and the other is 2x+y=-16

OpenStudy (anonymous):

no the question reads solve the following system of equations/ 3y=12 2x+y=-16 my answers include (-4,-10) (-10,4) (4, 10) (3,3)

OpenStudy (anonymous):

yes a system of equations is two equations. there is a space between the 12 and the 2x.

OpenStudy (anonymous):

can u help me solve this

OpenStudy (anonymous):

yes. they are asking to solve for x and y

OpenStudy (anonymous):

so start by solving for y 3y=12 y=12÷3 y=4

OpenStudy (anonymous):

now sub 4 in for y in the other equation to find x

OpenStudy (anonymous):

Would it be 4,10

OpenStudy (anonymous):

-10,4

OpenStudy (anonymous):

2x + y = -16 2x + 4 = -16 2x = -16 - 4 2x = -20 x = -20/2 x = -10 Its a - 10.